OpenStudy (anonymous):

integrate 2/(x^2 -2x) dx from x=3 to infinity

myininaya (myininaya):

\[\frac{2}{x^2-2x}=\frac{2}{x(x-2)}=\frac{A}{x}+\frac{B}{x-2}\] So you find A and B so that is true... What did you get for A and B?